### v1

# A discrete Benamou-Brenier formulation of Optimal Transport on graphs

Abstract
We propose a discrete transport equation on graphs which connects distributions on both vertices and edges. We then derive a discrete analogue of the Benamou-Brenier formulation for Wasserstein-$1$ distance on a graph and as a result classify all $W_{1}$ geodesics on graphs.
I Introduction
The classical transport problem considers two measures $\mu$ and $\nu$ on spaces $\mathcal{X}$ and $\mathcal{Y}$, where we have cost (and in our case a metric) of transporting $x\in\mathcal{X}$ to $y\in\mathcal{Y}$, which we denote $d(x,y)$. Then we ask: how can we transport $\mu$ to $\nu$ while minimising the expected cost? This gives us a measure of discrepancy between $\mu$ and $\nu$ relative to $d(x,y)$. The Kantorovich formulation of Optimal Transport [11] considers the minimum transport cost over couplings of $\mu$ and $\nu$.
Definition 1 (Kantorovich Formulation of optimal transport).
Let $\mu,\nu$ be a measures on $\mathcal{X}$ and $d$ a metric on $\mathcal{X}$, we define the set of couplings of $\mu$ and $\nu$ to be $\Pi(\mu,\nu):=\{\pi\in\mathcal{P}(\mathcal{X}^{2}):\pi(A\times\mathcal{X})=\mu(A),\pi(\mathcal{X}\times B)=\nu(B)\}$. Then we define the Wasserstein $p$ distance by:
$W_{p}(\mu,\nu)^{q}=\inf_{\pi\in\Pi(\mu,\nu)}\left(\int_{\mathcal{X}^{2}}d(x,y)^{p}d\pi\right).$ (1)
It is known that $W_{p}$ is a metric on the space of (finite $p$’th moment) measures on $\mathcal{X}$, see [20, Definition 6.4]. The Wasserstein distance is popular in machine learning circles, namely as a loss function [2, 8, 17] and utilising its Riemannian structure for Wasserstein Gradient Flows [10, 15]. One notable development connecting optimal transport and differential geometry came via the Benamou-Brenier formulation, which re-parametrised the minimisation in terms of time-dependent distribution and velocity field pairs $(f_{t},v_{t})$.
Proposition 1 (Benamou-Brenier Formulation).
Given two distributions $f(0),f(1)\in\mathcal{P}(\mathbb{R}^{d})$, we can then express
[…]
where $(f_{t},v_{t})$ are constrained by the transport equation - also known as the continuity equation:
$\partial_{t}f_{t}+\nabla\cdot(f_{t}v_{t})=0$ (3)
where $\nabla\cdot$ is the divergence operator, (see [5]).
Definition 2 (Constant Speed Geodesics).
Given a metric space $(\mathcal{X},d)$, we say a path $f_{t}$ is a Constant Speed Geodesic if and only if $d(f_{s},f_{t})=|s-t|d(f_{0},f_{1})$ and for a general path we define its Speed as $|\dot{f}|_{d}=\lim_{h\rightarrow 0}\frac{1}{h}d(f_{t+h},f_{t})$. Constant speed geodesics have constant speed, see [20, Remark 7.7].
The infimum in (2) is achieved by choosing $f_{t}$ as a constant speed $W_{p}$ geodesic and $v_{t}$ satisfying $v_{t}||=|\dot{f}|_{W_{p}}$, (see [5]).
On $\mathbb{R}^{d}$, the constant speed $W_{p}$ geodesics are the McCann displacement interpolations, see [14, Definition 1.1]. Though for discrete measures, say on $\{0,1...,n\}\subset\mathbb{R}$, the McCann interpolations will traverse $\mathbb{R}$, but not remain in the domain $\{0,1...,n\}$. This means to interpolate across discrete domains, considering them as embedded in a continuum will not suffice. This unfortunately sacrifices the classical Benamou-Brenier formula.
Work has already been done to amend this. Maas [13], gave a discrete transport equation and a new metric $\mathcal{W}$, distinct from $W_{2}$, which minimised a kinetic energy functional and interplayed well with discrete gradient flows.
Beckmann [4] provided a (time independent) velocity field formulation of $W_{1}$ on both continuous and discrete domains which preceded Benamou and Brenier. A concise reference can be found in [18, Section 6.3].
Finally, Hillion and Johnson [9] provided conditions for a discrete Benamou-Brenier formulation on $\mathbb{Z}$, which we are building on. In this paper we offer our own discrete transport equation including triples $(f,v,g)$, and derive a Benamou-Brenier formulation for $W_{1}$ on trees (Lemma 2, Theorem 1). We then generalise this result to graphs (Lemma 3, Theorem 2) and classify the constant speed $W_{1}$ geodesics via the solutions to our formulation (Proposition 4).
II Notation and the Discrete Transport Equation
Example 1.
Consider the path of Binomial distributions $f(t)=\text{Bin}(n,p(t))$ on the set $\mathcal{V}=\{0,1,...,n\}$ and $p$ varies with time. We show the case $n=5$ below:
$f_{0}$$f_{1}$$f_{2}$$f_{3}$$f_{4}$$f_{5}$
By the product rule and some rudimentary combinatorics we can express the derivative as:
$\partial_{t}f_{x}$ $=np^{\prime}\text{Bin}_{x-1}(n-1,p)-np^{\prime}\text{Bin}_{x}(n-1,p)$
$=-\nabla_{1}(np^{\prime}\text{Bin}_{x}(n-1,p))$
where $\nabla_{1}(h_{x}):=h_{x}-h_{x-1}$ is the finite difference operator.
This resembles (3), yet the distribution inside the spatial derivative is no longer $f$, in fact it is another distribution $g$, that has support $\{0,1,...,n-1\}$. This means the discrete transport equation satisfied in this case is
$\partial_{t}f+\nabla_{1}(vg)=0$
where $v(t)=np^{\prime}$ and $g(t)=\text{Bin}(n-1,p)$. Our key insight is that $\mathcal{V}$ can be considered as the vertices of a path graph $\mathcal{G}$, and that $g$ and $v$ exists on the edges of $\mathcal{G}$ as follows:
$f_{0}$$f_{1}$$f_{2}$$f_{3}$$f_{4}$$f_{5}$$v_{0}g_{0}$$v_{1}g_{1}$$v_{2}g_{2}$$v_{3}g_{3}$$v_{4}g_{4}$
and the discrete transport equation states that the rate of change at node $x$ is the potential difference across that node. On a general graph, we do not have the canonical rightward direction of edges for mass to flow, so in some sense on a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ we must choose this orientation arbitrarily.
II-A Matrices, Operators, Equations and Integrals
In this section, we will define our necessary objects. This includes a divergence operator, a discrete transport equation, and an energy functional, which we aim to minimise, as constrained by the discrete transport equation.
Definition 3 (Incidence Matrix and Arrow Shorthand).
Given a directed graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, we define the incidence matrix $\Omega=(\omega_{x,k})_{x\in\mathcal{V},k\in\mathcal{E}}$, by
[…]
As shorthand for $\omega_{x,k}=1$ and $\omega_{k,x}=-1$ we will write $x\rightarrow k$ and $k\rightarrow x$ respectively.
For example, we could orient $\mathbb{Z}$ by directing edges away from 0, rather than left to right.
Definition 4 (Floor and Ceiling Notation).
Given an edge $k\in\mathcal{E}$, we denote $\lceil k\rceil$ to be the node it is incoming to, and $\lfloor k\rfloor$ to be the node it is outgoing from.
Now, we can use the incidence matrix to define discrete derivative and divergence operators between functions on vertices and edges respectively.
Definition 5 (The Gradient and Divergence Operator).
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be a graph with incidence matrix $\Omega$. Then
  1. 1.
    For a function $f$ defined on vertices $\mathcal{V}$, we define the Gradient Operator at edge $k\in\mathcal{E}$ to be:
    $(\nabla f)_{k}:=(-\Omega^{T}\cdot f)_{k}=-\sum_{x\in\mathcal{V}}\omega_{x,k}f_{x}=f_{\lfloor k\rfloor}-f_{\lceil k\rceil},$
    i.e., the difference in function $f$ along the edge $k$.
  2. 2.
    For a function $g$ defined on the edges $\mathcal{E}$, we define the Divergence Operator at vertex $x\in\mathcal{V}$ to be:
    $(\nabla\cdot g)_{x}:=-\Omega\cdot g=-\sum_{k\in\mathcal{E}}\omega_{x,k}g_{k}=\sum_{k:x\rightarrow k}g_{k}-\sum_{k:k\rightarrow x}g_{k},$
    i.e., the difference in total flow out of and flow into $x$.
If we consider $\nabla_{1}$ as instead defined on edges (x,x1)(x,x-1) of $\mathbb{Z}$, it corresponds exactly to the $\nabla\cdot$ operator. The composition $\mathcal{L}\coloneqq\Omega\Omega^{T}$ is the Laplacian.
II-B The Discrete Transport Equation
We now build a framework to generalise Example 1. We start by building the transport equation as satisfied by a triple $(f,v,g)$, where we impose that $v$ and $g$ exist on edges.
Definition 6 (Discrete Transport Equation).
Given a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ with incidence matrix $\Omega$, we say the transport equation is satisfied if: for a path of distributions on vertices $f=f(t)_{x}$, a path of functions on edges $v=v(t)_{k}$ (called the velocity), and a path of distributions on edges $g=g(t)_{k}$, we have
$\partial_{t}f(t)_{x}+(\nabla\cdot v(t)g(t))_{x}\text{ for all }x\in\mathcal{V},t\in[0,1].$ (4)
This can be compactly expressed in terms of the incidence matrix $\Omega$:
$\partial_{t}f=\Omega\cdot(vg)$ (5)
where the product $vg$ is component-wise.
Example 2.
Consider the following directed star graph $S_{3}$;
[…]
and define $f$ to be $f(t)=Z(1,s,s,s)$ for an arbitrary decreasing and positive $s(t)$ and partition function $Z(t)=\frac{1}{1+3s}$. This is the stationary distribution of a Markov jump process $Q$ where $q_{x0}s(t)=q_{0x}$. Equation (4) states that
$\partial_{t}f_{0}$ $=-Z^{\prime}(t)$ $=-v_{e_{1}}g_{e_{1}}-v_{e_{2}}g_{e_{2}}-v_{e_{3}}g_{e_{3}},$
$\partial_{t}f_{1}$ $=\frac{1}{3}Z^{\prime}(t)$ $=v_{e_{1}}g_{e_{1}},$
$\partial_{t}f_{2}$ $=\frac{1}{3}Z^{\prime}(t)$ $=v_{e_{2}}g_{e_{2}},$
$\partial_{t}f_{3}$ $=\frac{1}{3}Z^{\prime}(t)$ $=v_{e_{3}}g_{e_{3}}.$
From here we have many choices for $(v,g)$. One of note is when $v$ is invariant across edges, i.e., when we choose
$v_{e_{1}}=v_{e_{2}}=v_{e_{3}}=Z^{\prime}(t),\quad g_{e_{1}}=g_{e_{2}}=g_{e_{3}}=\frac{1}{3}.$
We will return to this distribution and will motivate our choice of solution in Example 5.
We have our discrete transport equation, but a Benamou-Brenier formulation also requires an energy functional to minimise. In Proposition 1, we compute the expectation of $v$ with respect to $f$. However since $v$ exists on the edges with $g$, it is more natural to take the expectation over $g$. We keep the integral over time unchanged.
Definition 7 (Integral Formulation).
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be a graph with incidence matrix $\Omega$ and suppose $(f,v,g)$ satisfy (4). For $q\geq 1$ we define
$\mathcal{I}_{q}(v,g)\coloneqq\left(\int_{0}^{1}\sum_{k\in\mathcal{E}}g_{k}(t)|v_{k}(t)|^{q}dt\right)^{\frac{1}{q}}.$ (6)
Then for two distributions $f(0),f(1)$ on $\mathcal{V}$, we define
$V_{q}(f(0),f(1))\coloneqq\inf_{f,v,g}\left\{\mathcal{I}_{q}(v,g):\partial_{t}f=\Omega\cdot(vg)\right\}$ (7)
and so $V_{q}$ is a measure of discrepancy between $f(0)$ and $f(1)$. This is a generalisation of a functional in [9, Definition 3.2], which is the work we are building on.
III Main Results
III-A Benamou-Brenier on the Vertices of a Tree
The Wasserstein-$1$ distance on $\mathbb{Z}$ has a closed form expression, see [6, Lemma 8.2]. Slightly lesser known is that a generalised expression for the tree exists too – thanks to Evans, [7] – which remarkably corresponds to a metric from Microbiology called the UniFrac metric, see [12]. Evans defines $W_{1}$ in terms of cuts, but for our purposes, we will formulate it in terms of tails. Let $\mathcal{G}$ be a tree, and choose an arbitrary root node $r\in\mathcal{V}$ and define our incidence matrix by directing arrows away from $r$. This allows us to define tails.
Definition 8 (Tails of Distributions).
We define the tail past $x\in\mathcal{V}$ to be
$U(x)=\{y\in\mathcal{V}:\exists k_{1},...k_{n}(x\rightarrow k_{1}\rightarrow\cdots k_{n}\rightarrow y)\}$
and then $F_{x}=\sum_{y\in\mathcal{V}}\mathbb{I}(y\in U(x))f_{y}$ is called the tail distribution of $f$ past $x$.
$U(x)$ and $F_{x}$ generalise the set $\{x,x+1,x+2,$...\} for $\mathbb{Z}$ and the tail distribution on $\mathbb{Z}$ respectively. Then we have the following result.
Proposition 2 (Wasserstein-$1$ on a Tree [7]).
Given a rooted tree $\mathcal{G}$, for two distributions $f(0)$ and $f(1)$ on $\mathcal{V}$ we have that
$W_{1}(f(0),f(1))=\sum_{x\in\mathcal{V}}|F(1)_{x}-F(0)_{x}|.$ (8)
It is worth noting that we have other special cases of $W_{1}$. For cyclic graphs, we have a another closed form expression which utilise cuts, see [19] and for a general graph, $W_{1}$ can be expressed as the minimum transport cost across possible spanning trees of $\mathcal{G}$, see [16].
Lemma 1 (Tail Variant of the Transport Equation).
Let $\mathcal{G}$ be a tree with root node $r$ and incidence matrix $\Omega$, if we denote the tail of $f$ by $F(t)_{x}\coloneqq\sum_{y}\mathbb{I}(y\in U(x))f(t)_{y}$, we can invert (4) via the tail distributions:
$\partial_{t}F(t)_{\lceil k\rceil}=v(t)_{k}g(t)_{k},$ (9)
and $g$ can be derived exactly from $f$ and $v$.
Proof.
We drop the dependence on $t$ for now, and consider:
$\partial_{t}F_{\lceil k\rceil}$ $=\sum_{y\in\mathcal{V}}\mathbb{I}(y\in U(\lceil k\rceil))\partial_{t}f_{y}$
$=\sum_{y\in\mathcal{V}}\mathbb{I}(y\in U(\lceil k\rceil))\sum_{j\in\mathcal{E}}\omega_{y,j}v_{j}g_{j}$
$=\sum_{j\in\mathcal{E}}v_{j}g_{j}\sum_{y\in\mathcal{V}}\omega_{y,j}\mathbb{I}(y\in U(\lceil k\rceil))$
$=\sum_{j\in\mathcal{E}}v_{j}g_{j}(\nabla\mathbb{I}(y\in U(\lceil k\rceil))_{j}$
$=\sum_{j\in\mathcal{E}}v_{j}g_{j}\mathbb{I}(j=k)$
$=v_{k}g_{k}.$
Then $g$ can be derived from $f$ and $v$ by rearranging (9). ∎
While we have inverted (4), we still have a family of possible $(v,g)$ which satisfy (9), below we discuss the most important case for our purposes.
Lemma 2 (Constant Speed Solutions).
Given a path $f$, we define the Constant Speed Solution to be
$v_{k}=\text{sign}(\partial_{t}F_{\lceil k\rceil})|v|,\quad g_{k}=|\partial_{t}F_{\lceil k\rceil}|/|v|$ (10)
where $|v|=\sum_{x\in\mathcal{V}}|\partial_{t}F(t)_{x}|$. For this we have $|v|=|\dot{f}|_{W_{1}}$, and if $f$ is a constant speed geodesic, we additionally have that $\mathcal{I}_{2}(v,g)=|v|=W_{1}(f(0),f(1))$.
Proof.
See Appendix -A. ∎
Theorem 1 (Benamou-Brenier across a tree).
Let $\mathcal{G}$ be a rooted tree, then we have
$W_{1}(f(0),f(1))=\inf_{f,g,v}\left\{\mathcal{I}_{2}(v,g):\partial_{t}=\Omega\cdot(vg))\right\}.$ (11)
for all $f(0)$ and $f(1)$ on $\mathcal{V}$. The minimum is achieved by a constant speed geodesic $f$, and the velocity $v$ satisfies $|v(t)_{k}|=|\dot{f}|_{W_{1}}$, analogous to the continuous case.
Proof.
An immediate consequence of (8) and (9) is we are given an integral formulation of $W_{1}$. Remarkably, if $(f,v,g)$ is any triple which satisfies (4), then we can write
$W_{1}(f(0),f(1)$ $=\sum_{x\in\mathcal{V}}|F(1)_{x}-F(0)_{x}|$
$=\sum_{x\in\mathcal{V}}\left|\int_{0}^{1}\partial_{t}F(t)_{x}dt\right|$
$=\sum_{k\in\mathcal{E}}\left|\int_{0}^{1}g(t)_{k}v(t)_{k}dt\right|$
which transforms a sum over tails to a sum over edges. Squaring both sides and applying both Jensen’s and the Cauchy Schwarz inequalities bounds $W_{1}$ above by $V_{2}$.
$W_{1}(f(0),f(1))^{2}$ $=\left(\sum_{k\in\mathcal{E}}\left|\int_{0}^{1}g(t)_{k}v(t)_{k}dt\right|\right)^{2}$
$\leq\left(\int_{0}^{1}\sum_{k\in\mathcal{E}}|g(t)_{k}v(t)_{k}|dt\right)^{2}$
$\leq\int_{0}^{1}\left(\sum_{k\in\mathcal{E}}|g(t)_{k}v(t)_{k}|\right)^{2}dt$
$\leq\int_{0}^{1}\sum_{k\in\mathcal{E}}g(t)_{k}|v(t)_{k}|^{2}dt=\mathcal{I}_{2}(v,g)^{2}$
So, we have that $V_{2}\geq W_{1}$ too. However, Lemma 2, says that a constant speed solution achieves $W_{1}$, so choosing an arbitrary constant speed geodesic as in Definition 2 - see Appendix -B for an explicit example - we have that $V_{2}(f(0),f(1))=W_{1}(f(0),f(1))$ and we are done. ∎
In fact, the proofs for Lemma 2 and Theorem 1 work for $q\geq 1$ too, so we have a more general result that $V_{q}=W_{1}$, (see Appendix. -C) and we can choose $q$ arbitrarily.
Example 3.
Consider Example 1 when p(t)=(1t)p+tqp(t)=(1-t)p+tq for $p>q>0$. We know that v(t)=n(qp)v(t)=n(q-p) and $g(t)=\text{Bin}(n-1,p(t))$ satisfy (4). We also have that $\mathcal{I}_{2}(v,g)=n(q-p)=W_{1}(f(0),f(1))$, so the original choice of $(f,v,g)$ is minimizing. This solution also contains a constant speed geodesic in the form of $f$.
Example 4.
Let $f(t)=\text{Poi}((1-t)\lambda+t\mu)$, then $v(t)=\mu-\lambda$ and $g(t)=\text{Poi}((1-t)\lambda+t\mu)$ also satisfy (4) and $\mathcal{I}_{2}(v,g)=\lambda-\mu=W_{1}(f(0),f(1))$. Again $f$ is a constant speed geodesic.
Example 5.
Consider Example 2. When $s(t)=\frac{-1}{3}(1+\frac{1}{at+b})$ for a,b<1a,b<-1, we have that $s$ is both decreasing and positive. We can also show that $Z^{\prime\prime}(t)=0$. So by choosing the edge invariant solution we know $Z^{\prime}(t)=v_{01}=v_{02}=v_{03}$ are also constant in time. Plugging this into $\mathcal{I}_{2}$ we have
$\mathcal{I}_{2}(v,g)=\int_{0}^{1}Z^{\prime}(t)dt=Z(1)-Z(1)=W_{1}(f(0),f(1)).$
For this choice of $s(t)$, we have that f(t)=(1t)f(0)+tf(1)f(t)=(1-t)f(0)+tf(1) and by Appendix -B this is a constant speed geodesic.
We cannot apply the same argument to the graph case as in the tree, to work toward the graph, we will build a Reduced Formulation of $V_{q}$ by studying the solutions to (4) in more detail.
III-B The Reduced Formulation on a Graph
In the continuous formulation, so long as smoothness assumptions were made, switching from $f_{t}$ to $v_{t}$ and vice versa was a matter of solving a first-order PDE, see [1, Chapter 16.1] for details. However, in our case, we can solve this more directly - although different methods are required for each direction.
Remark 1.
Given $f$, solving (5) for $(v,g)$ involves inverting $\Omega$. Fortunately there is a plethora of work on solving these systems, see [3, Section 2]. For an incidence matrix $\Omega$, we have $\text{Im}(\Omega)=\{x:\sum_{j}x_{j}=0\}$ and $\operatorname{rank}(\Omega)=|\mathcal{V}|-1$, see [3, Lemma 2.4]. We know $\partial_{t}f\in\text{Im}(\Omega)$ for all $t$, so we can always solve for the vector $vg$. Although, this system is not full rank. Removing an arbitrary row $r$ from $f$ and $\Omega$ gives
$\partial_{t}\tilde{f}=\tilde{\Omega}\cdot(vg)$
which is full rank. Then for an arbitrary right inverse $P$ (such that $\tilde{\Omega}P=I$) we can express the general solution as
$vg=P\cdot\partial_{t}\tilde{f}+\epsilon$
for some $\epsilon\in\ker(\Omega)$. One can find a right inverse by choosing a spanning tree $\mathcal{T}$ and considering its path matrix $P_{\mathcal{T}}$, the right inverse then has block structure including $P_{\mathcal{T}}$. This additionally characterises $\ker(\Omega)$ as the cycle space of $\Omega$, see [3, Theorem 2.13]. In the case that $\mathcal{G}$ is a tree, $\tilde{\Omega}$ is invertible, and $\tilde{\Omega}^{-1}$ is defined by (9). For each valid $(P,\epsilon)$ we have a family of solutions for $(v,g)$.
Remark 2.
Suppose we instead have a pair $(v,g)$ on the edges of a graph, we can create a path $f$ on the vertices as follows:
$f(t)_{x}=\int_{0}^{t}\sum_{k\in\mathcal{E}}\omega_{x,k}v(\tau)_{k}g(\tau)_{k}d\tau+f(0)_{x}$ (12)
and for shorthand we denote $f=\int\Omega\cdot(vg)$. Conversely, for any path $f(t)$ on the vertices, no matter which $(v,g)$ pair we choose as a solution to (4) we can reconstruct $f$ by using (12).
The lack of a unique inverse prevents us from obtaining $g$ in terms of $f$ and $v$ or $v$ in terms of $f$ and $g$. However we can use Remark 2, to limit our search for pairs $(v,g)$ living on the edges of $\mathcal{G}$.
Proposition 3 (Reduced Formulation).
We have that
$V_{q}(f(0),f(1))=\inf_{v,g}\left\{\mathcal{I}_{q}(v,g):f(1)-f(0)=\Omega\textstyle\int_{0}^{1}gvdt\right\}$
Proof.
By Remark 2, we have a surjection $(v,g)\mapsto f$ such that $\partial_{t}f=\Omega\cdot(vg)$, so searching over triples $(f,v,g)$ accounts to searching over $(v,g)$ and inducing a path $f$ via Equation (12). However we require the constructed $f$ to truly interpolate between $f(0)$ and $f(1)$, which can only happen if
$\int_{0}^{1}\sum_{k\in\mathcal{E}}\omega_{x,k}g(t)_{k}v(t)_{k}dt=f(1)-f(0).$ (13)
As shorthand we denote $\Omega\cdot\int_{0}^{1}gvdt=f(1)-f(0)$. ∎
This formulation is connected to the Beckmann formulation [4], albeit with time-dependent parameters.
Remark 3.
We have a new condition given by (13). So we should also investigate the structure of solutions, fortunately we can inherit most of the machinery from Remark 1, as we are still inverting an incidence matrix. We have that $(v,g)$ must satisfy
$\int_{0}^{1}vgdt=P\cdot[\tilde{f}(1)-\tilde{f}(0)]+\epsilon$ (14)
for some $P$ such that $\Omega\cdot P=I$ and $\epsilon\in\ker\Omega$.
III-C Benamou-Brenier on the Vertices of a graph
The multiple constant speed solutions explains why we cannot extend the argument of Section III-A to a graph. In fact not all constant speed solutions will minimize $V_{q}$, never mind integrate to $W_{1}$. However we can demonstrate that these solutions minimize $\mathcal{I}_{q}$ for a given $P$ and $\epsilon$.
Lemma 3 (Generalised Constant Speed Solutions).
Given a graph $\mathcal{G}$ and $(P,\epsilon)$ a valid choice for inverting $\Omega$, define the Constant Speed Solution associated to $(P,\epsilon)$ to be
$v_{k}=\text{sign}(P\cdot\tilde{f}+\epsilon)_{k}|v|,\quad g_{k}=|P\cdot\tilde{f}+\epsilon|_{k}/|v|$ (15)
where $|v|=\sum_{k\in\mathcal{V}}|P\cdot\tilde{f}+\epsilon|_{k}$. Then the constant speed solution minimizes $\mathcal{I}_{q}(v,g)$ over the pairs $(v,g)$ associated to $(P,\epsilon)$, i.e pairs which satisfy (14).
Proof.
See Appendix -D. ∎
On a tree, we have one inverse (up to choice of root node) and Lemma 3 is equivalent to the constant speed solution minimising $\mathcal{I}_{q}(v,g)$, which is implied by Theorem 1.
We now present the main result of the paper: showing that the Benamou-Brenier formula holds for $q\geq 1$ on a general connected graph.
Theorem 2 (Benamou-Brenier across a Graph).
For a connected graph $\mathcal{G}$, for any $q\geq 1$, we can express the Wasserstein-$1$ distance between $f(0)$ and $f(1)$ by
$\inf_{f,v,g}\left\{\left(\int_{0}^{1}\sum_{k\in\mathcal{E}}g(t)_{k}|v(t)_{k}|^{q}dt\right)^{\frac{1}{q}}:\partial_{t}f=\Omega\cdot(vg)\right\}$ (16)
where $\Omega$ is an arbitrary incidence matrix of $\mathcal{G}$, and the infimum is achieved for at least one triple $(f,v,g)$.
Proof.
See Appendix -E. ∎
Remark 4 (Wasserstein Tree Distance).
The proof of Theorem 2 is independent of the proof for the tree, meaning that the $W_{1}$ distance on a tree can be derived from Theorems 1 and 2.
The following proposition demonstrates our motivation for considering time dependent pairs $(v,g)$, instead of fixed flows along edges, in that they induce a family of constant speed $W_{1}$, geodesics.
Proposition 4 (Geodesics).
Given a pair $(v,g)$ which achieves $W_{1}$, the induced path $f$ is a constant speed $W_{1}$ geodesic. Conversely, if $f$ is a constant speed geodesic, then there exists $(v,g)$ which achieves $W_{1}$ and induces $f$.
Proof.
Suppose that $(v,g)$ minimises $\mathcal{I}_{1}$. If it has constant speed, proceed, else substitute it for a constant speed solution as in Lemma 2. Then we define the path $f(t)=\int\Omega(vg)d\tau$ for $t\in[0,1]$. For an arbitrary subinterval $[s,t]$, the triple $(f,v,g)$ restricted to $[s,t]$ is still a solution to (5). What is unclear, is whether it is a minimising solution too. Suppose there exists $\tilde{v},\tilde{g}$ such that
$\int_{s}^{t}\sum_{k\in\mathcal{E}}\tilde{g}(\tau)_{k}|\tilde{v}(\tau)_{k}|d\tau<\int_{s}^{t}\sum_{k\in\mathcal{E}}g(\tau)_{k}|v(t)_{k}|dt.$
[…]
$(\hat{v},\hat{g})=\begin{cases}(\tilde{v},\tilde{g})&:\tau\in[s,t]\\ (v,g)&:\tau\not\in[s,t].\end{cases}$
But then by linearity of the integral we will have $\mathcal{I}_{1}(\tilde{v},\tilde{g})<\mathcal{I}_{1}(v,g)$ and hence $(v,g)$ is not a minimiser. So $(v,g)$ is a minimising solution between $[s,t]$ too and we have
$W_{1}(f(s),f(t))$ $=\int_{s}^{t}|v|d\tau$
$=(t-s)\int_{0}^{1}|v|d\tau=|t-s|W_{1}(f(0),f(1))$
and so $f$ is a constant speed geodesic.
Now suppose $f$ is a constant speed geodesic, we will use the result that the $W_{1}$ distance on a graph $\mathcal{G}$ is equal to $W_{1}$ across one of the spanning trees of $\mathcal{G}$, which we call $\mathcal{T}$. Then, we choose $(v,g)$ as the constant speed solution as in Lemma 2. This pair achieves $W_{1}$ on $\mathcal{T}$ and hence on $\mathcal{G}$. ∎
Corollary 1 (Convex Interpolations are $W_{1}$ Geodesics).
Let $f(0),f(1)$ be two distributions on the vertices of a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, then f(t)=(1t)f(0)+tf(1)f(t)=(1-t)f(0)+tf(1) is a constant speed geodesic.
Proof.
Let $J$ be a minimising flow in the Beckmann formulation, then we choose, $v_{k}g_{k}=J_{k}$. This achieves $W_{1}$ and then $f(t)=\int_{0}^{t}\Omega\cdot Jd\tau=t(f(1)-f(0))+f(0)$ is a constant speed $W_{1}$ geodesic. ∎
IV Conclusion
We have extended the work of Hillion and Johnson (see [9, Lemma 3.6]) to demonstrate that the discrete Benamou-Brenier formulation of $W_{1}$ holds in generality on $\mathbb{Z}$, on trees, and eventually on graphs. The tree case lends itself particularly well, with an exact form for the minimizing triples in terms of tail distributions of a constant speed geodesics.
While the minimal Beckmann flow provides us with a convex interpolation (see Corollary 1), we see that on $\mathbb{Z}$ and even cycles we have minimizing triples $(f,v,g)$ with more exotic geodesics (see Examples 3,4 and Appendix -F). We formalise this characterisation of constant speed $W_{1}$ geodesics in Proposition 4.
-A Proof of Lemma 2 for $q\geq 1$
Proof.
Let us compute $|\dot{f}|_{W_{1}}$ directly, we have that
$|\dot{f}|_{W_{1}}$ $=\lim_{h\rightarrow 0}\frac{W_{1}(f(t+h),f(t))}{h}$
$=\lim_{h\rightarrow 0}\frac{1}{h}\sum_{x\in\mathcal{V}}|F(t+h)_{x}-F(t)_{x}|$
$=\sum_{x\in\mathcal{V}}\left|\lim_{h\rightarrow 0}\frac{F(t+h)_{x}-F(t)_{x}}{h}\right|$
$=\sum_{x\in\mathcal{V}}|\partial_{t}F(t)_{x}|=|v|(t).$
Now assume that $f$ is a constant speed geodesic, we know from Definition 2 that
$W_{1}(f(t+h),f(t))=|h|W_{1}(f(0),f(1)),$
re-arranging and taking limits we have
$|v|=|\dot{f}|_{W_{1}}=\lim_{h\rightarrow 0}\frac{W_{1}(f(t+h),f(t))}{|h|}=W_{1}(f(0),f(1)).$
So overall we have
$\mathcal{I}_{q}(v,g)^{q}$ $=\int_{0}^{1}\sum_{k\in\mathcal{E}}g(t)_{k}v(t)_{k}^{q}dt$
$=\int_{0}^{1}\sum_{k\in\mathcal{E}}g(t)_{k}|v|^{q}dt$
$=\int_{0}^{1}|v|^{q}dt$
$=|v|^{q}=W_{1}(f(0),f(1))^{q}$
and so $\mathcal{I}_{q}(v,g)=W_{1}(f(0),f(1))$. ∎
-B Proof that Convex Interpolation is a Geodesic on a Tree
Proof.
Consider a convex interpolation f(t)=(1t)f(0)+tf(1)f(t)=(1-t)f(0)+tf(1), then the tail distribution is similarly convex F(t)=(1t)F(0)+tF(1)F(t)=(1-t)F(0)+tF(1). Then we directly compute:
$W_{1}(f(s),f(t))$ $=\sum_{x\in\mathcal{V}}|F(s)_{x}-F(t)_{x}|$
$=\sum_{x\in\mathcal{V}}|(1-s)F(0)_{x}+sF(1)_{x}$
$\quad\quad\quad-(1-t)F(0)_{x}-tF(1)_{x}|$
$=\sum_{x\in\mathcal{V}}|(t-s)(F(1)_{x}-F(0)_{x})|$
$=|t-s|\sum_{x\in\mathcal{V}}|F(1)_{x}-F(0)_{x}|$
$=|s-t|W_{1}(f(0),f(1)).$
So $f$ is a constant speed geodesic. ∎
-C Proof of Theorem 1 for $q\geq 1$
Proof.
An immediate consequence of (8) and (9) is we are given an integral formulation of $W_{1}$. Remarkably, if $(f,v,g)$ is any triple which satisfies (4), then we can write
$W_{1}(f(0),f(1)$ $=\sum_{x\in\mathcal{V}}|F(1)_{x}-F(0)_{x}|$
$=\sum_{x\in\mathcal{V}}\left|\int_{0}^{1}\partial_{t}F(t)_{x}dt\right|$
$=\sum_{k\in\mathcal{E}}\left|\int_{0}^{1}g(t)_{k}v(t)_{k}dt\right|$
which transforms a sum over tails to a sum over edges. Although this goes further, by applying both Jensen’s and the Cauchy Schwarz inequalities, we can bound $W_{1}$ above by $V_{2}$.
$W_{1}(f(0),f(1))^{q}$ $=\left(\sum_{k\in\mathcal{E}}\left|\int_{0}^{1}g(t)_{k}v(t)_{k}dt\right|\right)^{q}$
$\leq\left(\int_{0}^{1}\sum_{k\in\mathcal{E}}|g(t)_{k}v(t)_{k}|dt\right)^{q}$
$\leq\int_{0}^{1}\left(\sum_{k\in\mathcal{E}}|g(t)_{k}v(t)_{k}|\right)^{q}dt$
$\leq\int_{0}^{1}\sum_{k\in\mathcal{E}}g(t)_{k}|v(t)_{k}|^{q}dt=\mathcal{I}_{q}(v,g)^{q}$
So we have that $V_{q}\geq W_{1}$ too. However, Lemma 2, says that a constant speed solution achieves $W_{1}$, so we in fact have that $V_{q}(f(0),f(1))=W_{1}(f(0),f(1))$ and we are done. ∎
-D Proof of Lemma 3
Proof.
[…]
$\mathcal{I}_{q}(v,g)^{q}$ $=\int_{0}^{1}\sum_{k}g(t)_{k}|v(t)_{k}|^{q}dt$
$\geq\left(\int_{0}^{1}\sum_{k}|g(t)_{k}v(t)_{k}|dt\right)^{q}$
$\geq\left(\sum_{k}|\int_{0}^{1}g(t)_{k}v(t)_{k}dt|\right)^{q}$
$=\left(\sum_{k}|P\cdot(f(1)-f(0))+\epsilon|\right)^{q}$
$=\int_{0}^{1}|\tilde{v}|^{q}dt$
$=\int_{0}^{1}\sum_{k\in\mathcal{E}}\tilde{g}(t)_{k}|\tilde{v}(t)_{k}|^{q}dt=\mathcal{I}_{q}(\tilde{v},\tilde{g})^{q}$
-E Proof of Theorem 2
Proof.
We first use the reduced form for $V_{q}$:
$V_{q}(f(0),f(1))=\inf_{v,g}\left\{\mathcal{I}_{q}(v,g):f(1)-f(0)=\Omega\textstyle\int_{0}^{1}gvdt\right\}$
We can express the Beckmann formulation of the $W_{1}$ distance as follows:
$W_{1}(f(0),f(1))=\inf_{J}\left\{\sum_{k\in\mathcal{E}}|J_{k}|:\Omega\cdot J=f(1)-f(0)\right\}$
We begin by showing that $V_{q}\geq W_{1}$, let $(v,g)$ satisfy
$\Omega\cdot\int_{0}^{1}gvdt=f(1)-f(0).$
We see that identifying $J\coloneqq\int_{0}^{1}gvdt$ gives a valid solution to Beckmann’s formulation. Then we have that
$\mathcal{I}^{q}(v,g)^{q}$ $=\int_{0}^{1}\sum_{k}g(t)_{k}|v(t)_{k}|^{q}dt$
$\geq\int_{0}^{1}\left(\sum_{k}|g(t)_{k}v(t)_{k}|\right)^{q}dt$
$\geq\left(\int_{0}^{1}\sum_{k}|g(t)_{k}v(t)_{k}|dt\right)^{q}$
$\geq\left(\sum_{k}\int_{0}^{1}|g(t)_{k}v(t)_{k}|dt\right)^{q}$
$\geq\left(\sum_{k}\left|\int_{0}^{1}g(t)_{k}v(t)_{k}dt\right|\right)^{q}$
$=\left(\sum_{k}|J_{k}|\right)^{q}$
$\geq W_{1}(f(0),f(1))^{q}$
So we have additionally that $V_{q}\geq W_{1}$. Conversely, choose $J$ a minimising solution (achieves $W_{1}$) to the Beckmann formulation, then we can choose $|v|=\sum_{k}|J_{k}|$, $v(t)_{k}=\text{sign}(J_{k})|v|$ and $g(t)_{k}=|J_{k}|/|v|$, which satisfies $g(t)_{k}v(t)_{k}=J_{k}$. Additionally this pair $(v,g)$ is constant in time, so we have that
$\Omega\cdot\int_{0}^{1}gvdt=\Omega\cdot(gv)=\Omega\cdot J=f(1)-f(0).$
So $(v,g)$ satisfy the condition for $V_{q}$, then we have that
$\left(\sum_{k}|J_{k}|\right)^{q}$ $=|v|^{q}$
$=\sum_{k}g(t)_{k}|v|^{q}$
$=\sum_{k}g(t)_{k}|v(t)_{k}|^{q}$
$=\int_{0}^{1}\sum_{k}g(t)_{k}|v(t)_{k}|^{q}dt$
$\geq V_{q}(f(0),f(1))^{q}$
So we additionally have $W_{1}\geq V_{q}$, so overall we have that for any $f(0)$ and $f(1)$, we have
$W_{1}(f(0),f(1))=V_{q}(f(0),f(1)).$
-F Supplementary Results
The following results are true by virtue of $V_{q}=W_{1}$, although contain useful insights nevertheless, including a proof that Wasserstein-$1$ is bounded below by the total variation.
Lemma 4.
$V_{q}(f(0),f(1))\geq\text{TV}(f(0),f(1))$ for all $q\geq 1$, where TV is the total variational distance.
Proof.
Consider the reduced form of $V_{q}$, and let $(v,g)$ satisfy
$\Omega\cdot\int_{0}^{1}vgdt=f(1)-f(0).$
Writing this out explicitly we have
$|f(1)_{x}-f(0)_{x}|$ $=\left|\int_{0}^{1}\sum_{k\rightarrow x}g(t)_{k}v(t)_{k}-\sum_{x\rightarrow k}g(t)_{k}v(t)_{k}dt\right|$
$\leq\int_{0}^{1}\left|\sum_{k\rightarrow x}g(t)_{k}v(t)_{k}-\sum_{x\rightarrow k}g(t)_{k}v(t)_{k}\right|dt$
$\leq\int_{0}^{1}\sum_{k\rightarrow x}|g(t)_{k}v(t)_{k}|+\sum_{x\rightarrow k}|g(t)_{k}v(t)_{k}|dt.$
But considering the identity
$\sum_{x}\sum_{k:k\rightarrow x}h_{k}+\sum_{x}\sum_{k:x\rightarrow k}h_{k}=2\sum_{k}h_{k}$
we have that
$\sum_{x}|f(1)_{x}-f(0)_{x}|$ $\leq 2\int_{0}^{1}\sum_{k}|g(t)_{k}v(t)_{k}|dt.$
Finally we have that
$\mathcal{I}_{q}(v,g)^{q}$ $\geq\int_{0}^{1}\sum_{k}g(t)_{k}|v(t)_{k}|^{q}dt$
$\geq\int_{0}^{1}\left(\sum_{k}|g(t)_{k}v(t)_{k}|\right)^{q}dt$
$\geq\left(\int_{0}^{1}\sum_{k}|g(t)_{k}v(t)_{k}|\right)^{q}$
$\geq\left(\frac{1}{2}\sum_{x}|f(1)_{x}-f(0)_{x}|\right)^{q}=\text{TV}(f(0),f(1))^{q}.$
So $V_{q}\geq\text{TV}$ too. ∎
Proposition 5.
$V_{q}$ defines a metric on $\mathcal{P}(\mathcal{V})$.
Proof.
We begin with positive definiteness, we know $\mathcal{I}_{q}(v,g)\geq 0$ for all $(v,g)$, so $V_{q}\geq 0$. Additionally if $V_{q}(f(0),f(1))=0$, then $\text{TV}(f(0),f(1))=0$ too, so $f(0)=f(1)$ since the total variation is a metric. Now we prove symmetry. For any pair $(v,g)$ which satisfies
$\Omega\cdot\int_{0}^{1}gvdt=f(1)-f(0)$
[…]
$\Omega\cdot\int_{0}^{1}\tilde{g}\tilde{v}dt$ $=-\Omega\cdot\int_{0}^{1}g(1-t)v(1-t)dt$
$=-\Omega\cdot\int_{0}^{1}g(t)v(t)dt$
$=f(0)-f(1).$
[…]
$V_{q}(f(0),f(1))$ $=\lim_{n\rightarrow\infty}\mathcal{I}_{q}(v^{(n)},g^{(n)})$
$=\lim_{n\rightarrow\infty}\mathcal{I}_{q}(\tilde{v}^{(n)},\tilde{g}^{(n)})$
$\geq V_{q}(f(1),f(0)).$
[…]
$\hat{v}(t),\hat{g}(t)=\begin{cases}\frac{1}{\rho}v\left(\frac{t}{\rho}\right),g\left(\frac{t}{\rho}\right)&:t\in[0,\rho]\\ \frac{1}{1-\rho}\tilde{v}\left(\frac{t-\rho}{1-\rho}\right),\tilde{g}\left(\frac{t-\rho}{1-\rho}\right)&:t\in[\rho,1]\end{cases}$
It can be shown that
$\mathcal{I}_{q}(\hat{v},\hat{g})^{q}=\frac{1}{\rho}\mathcal{I}_{q}(v,g)^{q}+\frac{1}{1-\rho}\mathcal{I}_{q}(\tilde{v},\tilde{g})^{q}$
and we choose $\rho$ such that
$\mathcal{I}_{q}(\hat{v},\hat{g})^{q}=\left(\mathcal{I}_{q}(v,g)+\mathcal{I}_{q}(\tilde{v},\tilde{g})\right)^{q}$
and hence
[…]
Then if we let $(g^{(n)},v^{(n)})_{n}$ and $(\tilde{g}^{(n)},\tilde{v}^{(n)})_{n}$ be sequences such that $\mathcal{I}_{q}(g^{(n)},v^{(n)})\rightarrow V_{q}(f(0),f^{*})$ and $\mathcal{I}_{q}(\tilde{g}^{(n)},\tilde{v}^{(n)})\rightarrow V_{q}(f^{*},f(1))$ respectively, then
$V_{q}(f(0),f(1))$ $\leq\lim_{n\rightarrow\infty}\mathcal{I}(\hat{v}^{(n)},\hat{g}^{(n)})$
$=V_{q}(f(0),f^{*})+V_{q}(f^{*},f(1))$
and we have the triangle inequality. Finally we motivate why we can choose $\rho$ in such a way. We are essentially choosing $\rho$ such that
$\frac{1}{\rho}a^{q}+\frac{1}{1-\rho}b^{q}=(a+b)^{q}.$
This is a quadratic in $\rho$ which has valid solutions in $[0,1]$. ∎
The following is a worked example for finding valid minimizing solutions on a $4$-cycle.
Example 6.
Consider the following distribution $f$ on a square generated by two probabilities $p$ and $q$, we consider the Wasserstein distance between when $p=p_{0},q=q_{0}$ and $p=p_{1},q=q_{1}$ respectively.
pqpqq(1p)q(1-p)(1p)(1q)(1-p)(1-$q)p(1$-q)0112233
Since these are product measures, we can show that $W_{1}(f(0),f(1))=|p_{1}-p_{0}|+|q_{1}-q_{0}|$. We can express the transport equation as:
$\partial_{t}f(t)_{0}$ $=p^{\prime}q+pq^{\prime}$ $=-g_{01}v_{01}-g_{03}v_{03}$
$\partial_{t}f(t)_{1}$ $=q^{\prime}(1-p)-p^{\prime}q$ $=\quad g_{01}v_{01}-g_{12}v_{12}$
$\partial_{t}f(t)_{2}$ $=-p^{\prime}(1-q)-q^{\prime}(1-p)$ $=\quad g_{12}v_{12}+g_{32}v_{32}$
$\partial_{t}f(t)_{3}$ $=p^{\prime}(1-q)-pq^{\prime}$ $=\quad g_{03}v_{03}-g_{32}v_{32}$
Since we have one cycle, we know $\text{Nullity}(\Omega)=1$, so we have one degree of freedom in solutions which are written as
$g_{01}v_{01}$ $=-p^{\prime}q+\epsilon$
$g_{03}v_{03}$ $=-pq^{\prime}-\epsilon$
$g_{12}v_{12}$ $=q^{\prime}(1-p)-\epsilon$
$g_{32}v_{32}$ $=p^{\prime}(1-q)+\epsilon$
We choose a specific constant speed solution $(v,g)$ by assigning $|v|=|p^{\prime}|+|q^{\prime}|$, then defining the velocity $v$ by:
$v_{01}=v_{32}=\text{sign}(p^{\prime})|v|,\quad v_{03}=v_{12}=\text{sign}(q^{\prime})|v|$
and defining the edge distribution $g$ by
$|v|g_{01}$ $=|p^{\prime}|q,$ $\quad|v|g_{32}$ $=|p^{\prime}|(1-q)$
$|v|g_{03}$ $=|q^{\prime}|p,$ $\quad|v|g_{12}$ $=|q^{\prime}|(1-p)$
When $p^{\prime}$ and $q^{\prime}$ are constant, $|v|$ is additionally constant in time – as well as edge-wise. This constant speed solution achieves $W_{1}$, so by Proposition 4, the induced path $f(t)$ is a geodesic. This corresponds to interpolating $p(t)=(1-t)p_{0}+tp_{1}$ and $q(t)=(1-t)q_{0}+tq_{1}$.

### v2

@@
 Abstract
-We propose a discrete transport equation on graphs which connects distributions on both vertices and edges.
-We then derive a discrete analogue of the Benamou-Brenier formulation for Wasserstein-$1$ distance on a graph and as a result classify all $W_{1}$ geodesics on graphs.
+We propose a discrete transport equation on graphs.
+We then derive a discrete analogue of the Benamou-Brenier formulation for Wasserstein-$1$ distance on a graph and as a result classify all constant speed $W_{1}$ geodesics on graphs.
 I Introduction
@@
 Definition 1 (Kantorovich Formulation of optimal transport).
-Let $\mu,\nu$ be a measures on $\mathcal{X}$ and $d$ a metric on $\mathcal{X}$, we define the set of couplings of $\mu$ and $\nu$ to be $\Pi(\mu,\nu):=\{\pi\in\mathcal{P}(\mathcal{X}^{2}):\pi(A\times\mathcal{X})=\mu(A),\pi(\mathcal{X}\times B)=\nu(B)\}$.
-Then we define the Wasserstein $p$ distance by:
-$W_{p}(\mu,\nu)^{q}=\inf_{\pi\in\Pi(\mu,\nu)}\left(\int_{\mathcal{X}^{2}}d(x,y)^{p}d\pi\right).$ (1)
+Let $\mu,\nu$ be a measures on $\mathcal{X}$ and $d$ a metric on $\mathcal{X}$, we define the set of couplings of $\mu$ and $\nu$ to be $\Pi(\mu,\nu):=\{\pi\in\mathcal{P}(\mathcal{X}^{2}):\pi(A\times\mathcal{X})=\mu(A),\pi(\mathcal{X}\times B)=\nu(B)\}$, and the Wasserstein $p$ distance by:
+$W_{p}(\mu,\nu)^{p}=\inf_{\pi\in\Pi(\mu,\nu)}\left(\int_{\mathcal{X}^{2}}d(x,y)^{p}d\pi\right).$ (1)
 It is known that $W_{p}$ is a metric on the space of (finite $p$’th moment) measures on $\mathcal{X}$, see [20, Definition 6.4].
 The Wasserstein distance is popular in machine learning circles, namely as a loss function [2, 8, 17] and utilising its Riemannian structure for Wasserstein Gradient Flows [10, 15].
-One notable development connecting optimal transport and differential geometry came via the Benamou-Brenier formulation, which re-parametrised the minimisation in terms of time-dependent distribution and velocity field pairs $(f_{t},v_{t})$.
+One notable development in optimal transport came via Benamou and Brenier [5], who re-parametrised the minimisation in terms of time-dependent distribution and velocity field pairs $(f_{t},v_{t})$.
 Proposition 1 (Benamou-Brenier Formulation).
@@
 […]
-where $(f_{t},v_{t})$ are constrained by the transport equation - also known as the continuity equation:
-$\partial_{t}f_{t}+\nabla\cdot(f_{t}v_{t})=0$ (3)
+where $(f_{t},v_{t})$ are constrained by the transport equation:
+$\partial_{t}f_{t}+\nabla\cdot(f_{t}v_{t})=0,$ (3)
 where $\nabla\cdot$ is the divergence operator, (see [5]).
@@
 Constant speed geodesics have constant speed, see [20, Remark 7.7].
-The infimum in (2) is achieved by choosing $f_{t}$ as a constant speed $W_{p}$ geodesic and $v_{t}$ satisfying $v_{t}||=|\dot{f}|_{W_{p}}$, (see [5]).
+The infimum in (2) can be achieved by a constant speed $W_{p}$ geodesic $f_{t}$, and $v_{t}$ satisfying $v_{t}||=|\dot{f}|_{W_{p}}$, (see [5]).
 On $\mathbb{R}^{d}$, the constant speed $W_{p}$ geodesics are the McCann displacement interpolations, see [14, Definition 1.1].
-Though for discrete measures, say on $\{0,1...,n\}\subset\mathbb{R}$, the McCann interpolations will traverse $\mathbb{R}$, but not remain in the domain $\{0,1...,n\}$.
+For discrete measures, say on $\{0,1...,n\}\subset\mathbb{R}$, the McCann interpolations will traverse $\mathbb{R}$, but not remain in the domain $\{0,1...,n\}$.
 This means to interpolate across discrete domains, considering them as embedded in a continuum will not suffice.
-This unfortunately sacrifices the classical Benamou-Brenier formula.
-Work has already been done to amend this.
-Maas [13], gave a discrete transport equation and a new metric $\mathcal{W}$, distinct from $W_{2}$, which minimised a kinetic energy functional and interplayed well with discrete gradient flows.
-Beckmann [4] provided a (time independent) velocity field formulation of $W_{1}$ on both continuous and discrete domains which preceded Benamou and Brenier.
-A concise reference can be found in [18, Section 6.3].
-Finally, Hillion and Johnson [9] provided conditions for a discrete Benamou-Brenier formulation on $\mathbb{Z}$, which we are building on.
-In this paper we offer our own discrete transport equation including triples $(f,v,g)$, and derive a Benamou-Brenier formulation for $W_{1}$ on trees (Lemma 2, Theorem 1).
+On a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, computing (1) is a $|\mathcal{V}|^{2}$-dimensional linear program.
+However, Beckmann [4] gave a time-independent velocity field formulation of $W_{1}$ which on a graph is $|\mathcal{E}|$-dimensional, (see [18, Section 6.3]).
+In fact the Benamou-Brenier formulation can be viewed as a generalisation of Beckmann’s work which introduces time parameters.
+The main contribution to a discrete Benamou-Brenier analogue came from Erbar and Maas [13], who formulated an analogue of $W_{2}$ on discrete domains by considering a transport equation on graphs and a kinetic energy functional much like in Proposition 1.
+This metric is Riemannian and interplays well with discrete gradient flows, although importantly it is not the $W_{2}$ distance.
+This left the problem of a Benamou-Brenier interpretation of $W_{p}$ on graphs open.
+Later, Hillion and Johnson [9] provided sufficient conditions for a true discrete Benamou-Brenier form of $W_{1}$ on $\mathbb{Z}$, which is the work we are building on.
+In this paper we offer our own discrete transport equation involving triples $(f,v,g)$, and derive a Benamou-Brenier formulation for $W_{1}$ on trees (Lemma 2, Theorem 1).
 We then generalise this result to graphs (Lemma 3, Theorem 2) and classify the constant speed $W_{1}$ geodesics via the solutions to our formulation (Proposition 4).
@@
 Example 1.
-Consider the path of Binomial distributions $f(t)=\text{Bin}(n,p(t))$ on the set $\mathcal{V}=\{0,1,...,n\}$ and $p$ varies with time. We show the case $n=5$ below:
+Consider the path of Binomial distributions $f(t)=\text{Bin}(n,p(t))$ on the set $\mathcal{V}=\{0,1,...,n\}$ where $p(t)$ varies with time. We show the case $n=5$ below:
 $f_{0}$$f_{1}$$f_{2}$$f_{3}$$f_{4}$$f_{5}$
@@
 where $v(t)=np^{\prime}$ and $g(t)=\text{Bin}(n-1,p)$.
-Our key insight is that $\mathcal{V}$ can be considered as the vertices of a path graph $\mathcal{G}$, and that $g$ and $v$ exists on the edges of $\mathcal{G}$ as follows:
+Our key insight is that $\mathcal{V}$ can be considered as the vertices of a path graph $\mathcal{G}$, and that $g$ and $v$ exist on the edges of $\mathcal{G}$ as follows:
 $f_{0}$$f_{1}$$f_{2}$$f_{3}$$f_{4}$$f_{5}$$v_{0}g_{0}$$v_{1}g_{1}$$v_{2}g_{2}$$v_{3}g_{3}$$v_{4}g_{4}$
@@
 […]
-As shorthand for $\omega_{x,k}=1$ and $\omega_{k,x}=-1$ we will write $x\rightarrow k$ and $k\rightarrow x$ respectively.
+As shorthand for $\omega_{x,k}=1$ and $\omega_{k,x}=-1$ we will write $k\rightarrow x$ and $x\rightarrow k$ respectively.
 For example, we could orient $\mathbb{Z}$ by directing edges away from 0, rather than left to right.
@@
 If we consider $\nabla_{1}$ as instead defined on edges (x,x1)(x,x-1) of $\mathbb{Z}$, it corresponds exactly to the $\nabla\cdot$ operator.
-The composition $\mathcal{L}\coloneqq\Omega\Omega^{T}$ is the Laplacian.
 II-B The Discrete Transport Equation
@@
 Definition 6 (Discrete Transport Equation).
-Given a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ with incidence matrix $\Omega$, we say the transport equation is satisfied if: for a path of distributions on vertices $f=f(t)_{x}$, a path of functions on edges $v=v(t)_{k}$ (called the velocity), and a path of distributions on edges $g=g(t)_{k}$, we have
-$\partial_{t}f(t)_{x}+(\nabla\cdot v(t)g(t))_{x}\text{ for all }x\in\mathcal{V},t\in[0,1].$ (4)
+Given a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ with incidence matrix $\Omega$, we say the transport equation is satisfied if: for a path of distributions on vertices $f=f(t)_{x}$, a path of functions on edges $v=v(t)_{k}$ (called the velocity), and a path of distributions on edges $g=g(t)_{k}$, we have that for each $x\in\mathcal{V}$
+$\partial_{t}f(t)_{x}+(\nabla\cdot v(t)g(t))_{x}\text{ for all }t\in[0,1].$ (4)
 This can be compactly expressed in terms of the incidence matrix $\Omega$:
@@
 […]
-and define $f$ to be $f(t)=Z(1,s,s,s)$ for an arbitrary decreasing and positive $s(t)$ and partition function $Z(t)=\frac{1}{1+3s}$.
+and define $f$ to be $f(t)=(Z,Zs,Zs,Zs)$ for an arbitrary positive $s(t)$ and partition function $Z(t)=\frac{1}{1+3s}$.
 This is the stationary distribution of a Markov jump process $Q$ where $q_{x0}s(t)=q_{0x}$.
@@
 It is worth noting that we have other special cases of $W_{1}$.
-For cyclic graphs, we have a another closed form expression which utilise cuts, see [19] and for a general graph, $W_{1}$ can be expressed as the minimum transport cost across possible spanning trees of $\mathcal{G}$, see [16].
+For cyclic graphs, we have a another closed form expression, (see [19]), which utilises cuts much like Evans [7].
+For a general graph, $W_{1}$ can be expressed as the minimum transport cost across possible spanning trees of $\mathcal{G}$, see [16].
 Lemma 1 (Tail Variant of the Transport Equation).
-Let $\mathcal{G}$ be a tree with root node $r$ and incidence matrix $\Omega$, if we denote the tail of $f$ by $F(t)_{x}\coloneqq\sum_{y}\mathbb{I}(y\in U(x))f(t)_{y}$, we can invert (4) via the tail distributions:
-$\partial_{t}F(t)_{\lceil k\rceil}=v(t)_{k}g(t)_{k},$ (9)
+Let $\mathcal{G}$ be a tree with root node $r$ and incidence matrix $\Omega$, we can invert (4) via the tail distributions:
+$\partial_{t}F(t)_{\lceil k\rceil}=v(t)_{k}g(t)_{k}\text{ for each }k\in\mathcal{E},$ (9)
 and $g$ can be derived exactly from $f$ and $v$.
@@
 Let $\mathcal{G}$ be a rooted tree, then we have
-$W_{1}(f(0),f(1))=\inf_{f,g,v}\left\{\mathcal{I}_{2}(v,g):\partial_{t}=\Omega\cdot(vg))\right\}.$ (11)
+$W_{1}(f(0),f(1))=\inf_{f,g,v}\left\{\mathcal{I}_{2}(v,g):\partial_{t}f=\Omega\cdot(vg))\right\}.$ (11)
 for all $f(0)$ and $f(1)$ on $\mathcal{V}$. The minimum is achieved by a constant speed geodesic $f$, and the velocity $v$ satisfies $|v(t)_{k}|=|\dot{f}|_{W_{1}}$, analogous to the continuous case.
@@
 $=\sum_{k\in\mathcal{E}}\left|\int_{0}^{1}g(t)_{k}v(t)_{k}dt\right|$
-which transforms a sum over tails to a sum over edges. Squaring both sides and applying both Jensen’s and the Cauchy Schwarz inequalities bounds $W_{1}$ above by $V_{2}$.
+which transforms a sum over tails to a sum over edges. Squaring both sides and applying both Jensen’s and the Cauchy–Schwarz inequalities bounds $W_{1}$ above by $V_{2}$.
 $W_{1}(f(0),f(1))^{2}$ $=\left(\sum_{k\in\mathcal{E}}\left|\int_{0}^{1}g(t)_{k}v(t)_{k}dt\right|\right)^{2}$
@@
 So, we have that $V_{2}\geq W_{1}$ too.
-However, Lemma 2, says that a constant speed solution achieves $W_{1}$, so choosing an arbitrary constant speed geodesic as in Definition 2 - see Appendix -B for an explicit example - we have that $V_{2}(f(0),f(1))=W_{1}(f(0),f(1))$ and we are done.
+However, Lemma 2 says that a constant speed solution achieves $W_{1}$, so choosing an arbitrary constant speed geodesic as in Definition 2 - see Appendix -B for an explicit example - we have that $V_{2}(f(0),f(1))=W_{1}(f(0),f(1))$ and we are done.
 ∎
-In fact, the proofs for Lemma 2 and Theorem 1 work for $q\geq 1$ too, so we have a more general result that $V_{q}=W_{1}$, (see Appendix. -C) and we can choose $q$ arbitrarily.
+In fact, the proofs for Lemma 2 and Theorem 1 work for $q\geq 1$ too, so we have a more general result that $V_{q}=W_{1}$, see Appendix -C , and we can choose $q$ arbitrarily.
 Example 3.
-Consider Example 1 when p(t)=(1t)p+tqp(t)=(1-t)p+tq for $p>q>0$.
+Consider the Binomial case (Example 1) when p(t)=(1t)p+tqp(t)=(1-t)p+tq for $p>q>0$.
 We know that v(t)=n(qp)v(t)=n(q-p) and $g(t)=\text{Bin}(n-1,p(t))$ satisfy (4).
@@
 We know $\partial_{t}f\in\text{Im}(\Omega)$ for all $t$, so we can always solve for the vector $vg$.
-Although, this system is not full rank.
-Removing an arbitrary row $r$ from $f$ and $\Omega$ gives
+Although this system is not full rank, removing an arbitrary row $r$ from $f$ and $\Omega$ gives
 $\partial_{t}\tilde{f}=\tilde{\Omega}\cdot(vg)$
@@
 $f(t)_{x}=\int_{0}^{t}\sum_{k\in\mathcal{E}}\omega_{x,k}v(\tau)_{k}g(\tau)_{k}d\tau+f(0)_{x}$ (12)
-and for shorthand we denote $f=\int\Omega\cdot(vg)$. Conversely, for any path $f(t)$ on the vertices, no matter which $(v,g)$ pair we choose as a solution to (4) we can reconstruct $f$ by using (12).
-The lack of a unique inverse prevents us from obtaining $g$ in terms of $f$ and $v$ or $v$ in terms of $f$ and $g$.
+and for shorthand we denote $f=\int\Omega\cdot(vg)$.
+By construction, $(f,v,g)$ will satisfy (4).
+Conversely, for any path $f(t)$ on the vertices, no matter which $(v,g)$ pair we choose as a solution to (4) we can reconstruct $f$ by using (12).
+The lack of a unique inverse of $\Omega$ prevents us from obtaining $g$ in terms of $f$ and $v$ or $v$ in terms of $f$ and $g$.
 However we can use Remark 2, to limit our search for pairs $(v,g)$ living on the edges of $\mathcal{G}$.
@@
 Remark 3.
-We have a new condition given by (13).
-So we should also investigate the structure of solutions, fortunately we can inherit most of the machinery from Remark 1, as we are still inverting an incidence matrix.
+We have a new condition given by (13), so we should also investigate the structure of solutions, fortunately we can inherit most of the machinery from Remark 1, as we are still inverting an incidence matrix.
 We have that $(v,g)$ must satisfy
@@
 III-C Benamou-Brenier on the Vertices of a graph
-The multiple constant speed solutions explains why we cannot extend the argument of Section III-A to a graph.
-In fact not all constant speed solutions will minimize $V_{q}$, never mind integrate to $W_{1}$.
-However we can demonstrate that these solutions minimize $\mathcal{I}_{q}$ for a given $P$ and $\epsilon$.
+The multiple possible inversions of $\Omega$ explain why we cannot extend the argument of Section III-A to a graph.
+In fact even constant speed solutions defined in Lemma 2 might not minimize $V_{q}$, never mind integrate to $W_{1}$.
+However we can demonstrate that constant speed solutions are still minimal.
 Lemma 3 (Generalised Constant Speed Solutions).
-Given a graph $\mathcal{G}$ and $(P,\epsilon)$ a valid choice for inverting $\Omega$, define the Constant Speed Solution associated to $(P,\epsilon)$ to be
-$v_{k}=\text{sign}(P\cdot\tilde{f}+\epsilon)_{k}|v|,\quad g_{k}=|P\cdot\tilde{f}+\epsilon|_{k}/|v|$ (15)
-where $|v|=\sum_{k\in\mathcal{V}}|P\cdot\tilde{f}+\epsilon|_{k}$.
-Then the constant speed solution minimizes $\mathcal{I}_{q}(v,g)$ over the pairs $(v,g)$ associated to $(P,\epsilon)$, i.e pairs which satisfy (14).
+Given a graph $\mathcal{G}$, we say a pair $(v,g)$ which satisfies (13) is a Constant Speed Solution if $|v_{k}|=|v|$ for some function $|v|:[0,1]\rightarrow\mathbb{R}$.
+For any pair $(v,g)$ we have a constant speed solution $(v^{*},g^{*})$ such that $\mathcal{I}_{q}(v,g)\geq\mathcal{I}_{q}(v^{*},g^{*})$.
+Additionally, we can find another constant speed solution $(v^{**},g^{**})$ such that $|v^{**}_{k}|=|v^{**}|=c$ for $c\in\mathbb{R}$ and $\mathcal{I}_{q}(v^{*},g^{*})\geq\mathcal{I}_{q}(v^{**},g^{**})$.
 Proof.
 See Appendix -D. ∎
-On a tree, we have one inverse (up to choice of root node) and Lemma 3 is equivalent to the constant speed solution minimising $\mathcal{I}_{q}(v,g)$, which is implied by Theorem 1.
+The key takeaway from Lemma 3 is that if we find any pair $(v,g)$ which achieves $W_{1}$, we can transform it via $(v^{*},g^{*})$ and $(v^{**},g^{**})$ to $W_{1}$ achieving pairs which have desirable properties.
+One of these properties will be key in the main result of Section IV.
+On a tree, we have one inverse (up to choice of root node) for Remark 3 and we know the constant speed solutions are explicitly defined via Lemma 2.
 We now present the main result of the paper: showing that the Benamou-Brenier formula holds for $q\geq 1$ on a general connected graph.
@@
 The proof of Theorem 2 is independent of the proof for the tree, meaning that the $W_{1}$ distance on a tree can be derived from Theorems 1 and 2.
-The following proposition demonstrates our motivation for considering time dependent pairs $(v,g)$, instead of fixed flows along edges, in that they induce a family of constant speed $W_{1}$, geodesics.
+IV Characterisation of Geodesics
+On $\mathbb{R}^{d}$, we can achieve the minima of (2) by a pair $(f,v)$ including a constant speed geodesic $f$.
+We also know that constant speed $W_{p}$-geodesics between $f(0)$ and $f(1)$ are the McCann displacement interpolations – which are induced by an optimal couplings between $f(0)$ and $f(1)$.
+In Examples 3, 4 and 5, we saw that the minimizing triple $(f,v,g)$ for (6) included a constant speed geodesic $f$ in much the same way.
+We also know in general that the path $f$ can be obtained from $(v,g)$ via (12) much like the McCann interpolations.
+We now establish our own characterisation of these geodesics.
 Proposition 4 (Geodesics).
-Given a pair $(v,g)$ which achieves $W_{1}$, the induced path $f$ is a constant speed $W_{1}$ geodesic.
-Conversely, if $f$ is a constant speed geodesic, then there exists $(v,g)$ which achieves $W_{1}$ and induces $f$.
+Given a pair $(v,g)$ which achieves $W_{1}$ in Proposition 3, and satisfies $|v_{k}|=|v|=c$, the induced path $f$ is a constant speed $W_{1}$ geodesic.
+Conversely, if $f$ is a constant speed geodesic, then there exists $(v,g)$ such that $|v_{k}|=|v|=c$ which achieves $W_{1}$ and induces $f$.
 Proof.
-Suppose that $(v,g)$ minimises $\mathcal{I}_{1}$.
-If it has constant speed, proceed, else substitute it for a constant speed solution as in Lemma 2.
-Then we define the path $f(t)=\int\Omega(vg)d\tau$ for $t\in[0,1]$.
-For an arbitrary subinterval $[s,t]$, the triple $(f,v,g)$ restricted to $[s,t]$ is still a solution to (5).
-What is unclear, is whether it is a minimising solution too.
-Suppose there exists $\tilde{v},\tilde{g}$ such that
-$\int_{s}^{t}\sum_{k\in\mathcal{E}}\tilde{g}(\tau)_{k}|\tilde{v}(\tau)_{k}|d\tau<\int_{s}^{t}\sum_{k\in\mathcal{E}}g(\tau)_{k}|v(t)_{k}|dt.$
-[…]
-$(\hat{v},\hat{g})=\begin{cases}(\tilde{v},\tilde{g})&:\tau\in[s,t]\\ (v,g)&:\tau\not\in[s,t].\end{cases}$
-But then by linearity of the integral we will have $\mathcal{I}_{1}(\tilde{v},\tilde{g})<\mathcal{I}_{1}(v,g)$ and hence $(v,g)$ is not a minimiser.
-So $(v,g)$ is a minimising solution between $[s,t]$ too and we have
-$W_{1}(f(s),f(t))$ $=\int_{s}^{t}|v|d\tau$
-$=(t-s)\int_{0}^{1}|v|d\tau=|t-s|W_{1}(f(0),f(1))$
+Let $(v,g)$ be a constant speed solution that minimises $I_{1}$ and $|v_{k}|=|v|=c$, then induce the path $f(t)$.
+For an arbitrary subinterval $[s,t]$, the pair $(v,g)$ restricted to $[s,t]$ is still a solution to (3).
+It can also be shown by contradiction that $(v,g)$ restricted to $[s,t]$ also minimises $I_{1}$ across $[s,t]$, (see Appendix -F).
+Since $(v,g)$ is constant speed, we have
+$W_{1}(f(s),f(t))=\int_{s}^{t}|v|d\tau$ $=(t-s)\int_{0}^{1}|v|d\tau$
+$=(t-s)W_{1}(f(0),f(1))$
 and so $f$ is a constant speed geodesic.
-Now suppose $f$ is a constant speed geodesic, we will use the result that the $W_{1}$ distance on a graph $\mathcal{G}$ is equal to $W_{1}$ across one of the spanning trees of $\mathcal{G}$, which we call $\mathcal{T}$.
-Then, we choose $(v,g)$ as the constant speed solution as in Lemma 2.
-This pair achieves $W_{1}$ on $\mathcal{T}$ and hence on $\mathcal{G}$.
+Conversely, suppose $f$ is a constant speed geodesic, so $|\dot{f}|_{W_{1}}=W_{1}(f(0),f(1))$.
+We then use the remarkable result that (if we let $W^{\mathcal{H}}_{1}$ denote the Wasserstein distance across the graph $\mathcal{H}$) there exists some spanning tree $\mathcal{T}\subset\mathcal{G}$ such that $W^{\mathcal{T}}_{1}(f(0),f(1))=W^{\mathcal{G}}_{1}(f(0),f(1))$, (see [16]).
+We can choose $(v,g)$ as the constant speed solution on $\mathcal{E}_{\mathcal{T}}$ (as in Lemma 2), and $v=g=0$ for any $\mathcal{E}\backslash\mathcal{E}_{\mathcal{T}}$.
+This pair achieves $W_{1}^{\mathcal{T}}=W_{1}^{\mathcal{G}}$ and since $(f,v,g)$ satisfy the transport equation, $f$ is induced by $(v,g)$.
 ∎
@@
 ∎
-IV Conclusion
+V Conclusion
 We have extended the work of Hillion and Johnson (see [9, Lemma 3.6]) to demonstrate that the discrete Benamou-Brenier formulation of $W_{1}$ holds in generality on $\mathbb{Z}$, on trees, and eventually on graphs.
-The tree case lends itself particularly well, with an exact form for the minimizing triples in terms of tail distributions of a constant speed geodesics.
-While the minimal Beckmann flow provides us with a convex interpolation (see Corollary 1), we see that on $\mathbb{Z}$ and even cycles we have minimizing triples $(f,v,g)$ with more exotic geodesics (see Examples 3,4 and Appendix -F).
-We formalise this characterisation of constant speed $W_{1}$ geodesics in Proposition 4.
+The tree case lends itself particularly well, with an exact form for the minimizing triples in terms of tail distributions of constant speed , (see Lemma 1 and Theorem 1).
+While the minimal Beckmann flow provides us with a convex interpolation (see Corollary 1), we see that on $\mathbb{Z}$ and even cycles we have minimizing triples $(f,v,g)$ with more exotic geodesics (see Examples 3, 4 and Example 6 in Appendix H).
+In fact the existence of Corollary 1 and Example 3 tell us that
+$f_{1}(t)$ $=\text{Bin}(n,(1-t)p+tq)$
+$f_{2}(t)$ $=(1-t)\text{Bin}(n,p)+t\text{Bin}(n,q)$
+are two valid constant speed geodesics between $\text{Bin}(n,p)$ and $\text{Bin}(n,q)$. We formalise this characterisation of $W_{1}$ geodesics in Proposition 4.
 -A Proof of Lemma 2 for $q\geq 1$
@@
 Proof.
-[…]
-$\mathcal{I}_{q}(v,g)^{q}$ $=\int_{0}^{1}\sum_{k}g(t)_{k}|v(t)_{k}|^{q}dt$
-$\geq\left(\int_{0}^{1}\sum_{k}|g(t)_{k}v(t)_{k}|dt\right)^{q}$
-$\geq\left(\sum_{k}|\int_{0}^{1}g(t)_{k}v(t)_{k}dt|\right)^{q}$
-$=\left(\sum_{k}|P\cdot(f(1)-f(0))+\epsilon|\right)^{q}$
-$=\int_{0}^{1}|\tilde{v}|^{q}dt$
-$=\int_{0}^{1}\sum_{k\in\mathcal{E}}\tilde{g}(t)_{k}|\tilde{v}(t)_{k}|^{q}dt=\mathcal{I}_{q}(\tilde{v},\tilde{g})^{q}$
+Let $(v,g)$ be a solution to (13), we induce a path $f$, and then since $(f,v,g)$ satisfy the transport equation, for some $(P,\epsilon)$ we have $vg=P\cdot\partial_{t}f+\epsilon$.
+Then define $(v^{*},g^{*})$ by
+$v^{*}_{k}$ $=\text{sign}(P\cdot\partial_{t}f+\epsilon)_{k}|v^{*}|$
+$g^{*}_{k}$ $=|P\cdot\partial_{t}f+\epsilon|_{k}/|v^{*}|$
+$|v^{*}|$ $=\sum_{k\in\mathcal{E}}|P\cdot\partial_{t}f+\epsilon|_{k}$
+then we know $\int_{0}^{1}v^{*}g^{*}dt=P\cdot(f(1)-f(0))+\epsilon$ and so $\int_{0}^{1}\Omega v^{*}g^{*}dt=f(1)-f(0)$. Finally we have
+$\mathcal{I}_{q}(v,g)^{q}$ $=\int_{0}^{1}\sum_{k\in\mathcal{E}}g_{k}|v_{k}|^{q}dt$
+$\geq\int_{0}^{1}\left(\sum_{k\in\mathcal{E}}|g_{k}v_{k}|\right)^{q}dt$
+$=\int_{0}^{1}\left(\sum_{k}|P\cdot\partial_{t}f+\epsilon|_{k}\right)^{q}dt$
+$=\int_{0}^{1}|v^{*}|^{q}dt$
+$=\int_{0}^{1}\sum_{k}g^{*}_{k}|v_{k}^{*}|^{q}dt=\mathcal{I}_{q}(v^{*},g^{*})^{q}.$
+Now consider an increasing function $\phi:[0,1]\rightarrow[0,1]$ such that $\phi(0)=0$ and $\phi(1)=1$ and the following pair:
+$v^{**}(t)_{k}=\phi^{\prime}(t)\cdot v^{*}(\phi(t))_{k},\quad g^{**}(t)_{k}=g^{*}(\phi(t))_{k}.$
+By integration by substitution we know
+$\int_{0}^{1}\Omega\cdot v^{**}g^{**}dt=\int_{0}^{1}\Omega\ \cdot v^{*}g^{*}dt=f(1)-f(0).$
+So $(v^{**},g^{**})$ is a valid solution. We consider when
+$\left|\phi^{\prime}(t)\frac{dv^{*}}{d\phi}\right|=\phi^{\prime}(t)\left|\frac{dv^{*}}{d\phi}\right|=c.$ (16)
+We denote $s(\phi)=\left|\frac{dv^{*}}{d\phi}\right|$ and $S(y)=\int_{0}^{y}s(y)dy$ then solving (16) we have $S(\phi(t))=ct+d$.
+We know that $c=S(\phi(1))-S(\phi(0))=S(1)-S(0)$ and $d=S(\phi(0))=S(0)$.
+So the choice of
+$\phi(t)\coloneqq S^{-1}[(S(1)-S(0))t+S(0)]$
+gives us $|v^{**}|=c=S(1)-S(0)=\int_{0}^{1}s(t)dt$.
+Note that $S^{-1}$ exists because $s(y)>0$ so $S$ is increasing.
+This parametrisation has time-independent speed and by the chain rule and Jensen’s inequality we have
+$\mathcal{I}_{q}(v^{**},g^{**})^{q}$ $=\int_{0}^{1}|v^{**}|^{q}dt$
+$=|v^{**}|^{q}$
+$=\left(\int_{0}^{1}|v^{**}|dt\right)^{q}$
+$=\left(\int_{0}^{1}\phi^{\prime}(t)|v^{*}(\phi(t))|dt\right)^{q}$
+$=\left(\int_{0}^{1}|v^{*}(s)|ds\right)^{q}$
+$\leq\int_{0}^{1}|v^{*}(s)|^{q}ds=\mathcal{I}_{q}(v^{*},g^{*})^{q}$
 -E Proof of Theorem 2
@@
 We see that identifying $J\coloneqq\int_{0}^{1}gvdt$ gives a valid solution to Beckmann’s formulation. Then we have that
-$\mathcal{I}^{q}(v,g)^{q}$ $=\int_{0}^{1}\sum_{k}g(t)_{k}|v(t)_{k}|^{q}dt$
+$\mathcal{I}_{q}(v,g)^{q}$ $=\int_{0}^{1}\sum_{k}g(t)_{k}|v(t)_{k}|^{q}dt$
 $\geq\int_{0}^{1}\left(\sum_{k}|g(t)_{k}v(t)_{k}|\right)^{q}dt$
@@
 $\geq W_{1}(f(0),f(1))^{q}$
-So we have additionally that $V_{q}\geq W_{1}$.
+So we have that $V_{q}\geq W_{1}$.
 Conversely, choose $J$ a minimising solution (achieves $W_{1}$) to the Beckmann formulation, then we can choose $|v|=\sum_{k}|J_{k}|$, $v(t)_{k}=\text{sign}(J_{k})|v|$ and $g(t)_{k}=|J_{k}|/|v|$, which satisfies $g(t)_{k}v(t)_{k}=J_{k}$.
@@
 So $(v,g)$ satisfy the condition for $V_{q}$, then we have that
-$\left(\sum_{k}|J_{k}|\right)^{q}$ $=|v|^{q}$
+$W_{1}(f(0),f(1))^{q}$ $=\left(\sum_{k}|J_{k}|\right)^{q}$
+$=|v|^{q}$
 $=\sum_{k}g(t)_{k}|v|^{q}$
@@
 $W_{1}(f(0),f(1))=V_{q}(f(0),f(1)).$
--F Supplementary Results
+-F Proof of Local Minimisation
+Proof.
+Suppose $(v,g)$ is a minimising pair for $I_{1}$ across $[0,1]$, i.e we have that
+$\mathcal{I}_{1}(v,g)=\int_{0}^{1}\sum_{k\in\mathcal{E}}v_{k}g_{k}dt$ $=\min_{(v_{0},g_{0})}\mathcal{I}_{1}(v_{0},g_{0})$
+$=W_{1}(f(0),f(1)).$
+Then we denote
+$\mathcal{I}_{q}^{[s,t]}(v,g)=\left(\int_{s}^{t}\sum_{k\in\mathcal{E}}g(t)_{k}|v(t)|_{k}^{q}dt\right)^{\frac{1}{q}}.$
+This implies that $\mathcal{I}_{1}^{[0,1]}=\mathcal{I}_{1}$. Suppose $(v,g)$ is not minimal across some $[s,t]$ for $\mathcal{I}_{1}$, then there exists some $(v^{*},g^{*})$ such that
+$\mathcal{I}_{1}^{[s,t]}(v^{*},g^{*})<\mathcal{I}_{1}^{[s,t]}(v,g).$
+From here we construct a new solution across $[0,1]$:
+$(v^{**}(\tau)_{k},g^{**}(\tau)_{k})=\begin{cases}(v^{*}(\tau)_{k},g^{*}(\tau)_{k})&:\tau\in[s,t]\\ (v(\tau)_{k},g(\tau)_{k})&:\tau\not\in[s,t]\end{cases}$
+Then by decomposing
+$\mathcal{I}_{1}^{[0,1]}=\mathcal{I}_{1}^{[0,s]}+\mathcal{I}_{1}^{[s,t]}+\mathcal{I}_{1}^{[t,1]}$
+we have that
+$\mathcal{I}_{1}^{[0,1]}(v^{**},g^{**})$ $<\mathcal{I}_{1}^{[0,s]}(v,g)+\mathcal{I}_{1}^{[s,t]}(v,g)+\mathcal{I}^{[t,1]}(v,g)$
+$=\mathcal{I}_{1}^{[0,1]}(v,g)$
+which contradicts the minimality of $(v,g)$. ∎
+-G Supplementary Results
 The following results are true by virtue of $V_{q}=W_{1}$, although contain useful insights nevertheless, including a proof that Wasserstein-$1$ is bounded below by the total variation.